OpenStudy (anonymous):

If 2xy+y^2=5, then dy/dx+?

OpenStudy (anonymous):

implicit differentiation ;)

OpenStudy (anonymous):

yes help please

OpenStudy (anonymous):

2y + (dy/dx)(2x) + 2y(dy/dx) = 5. simplify.

OpenStudy (anonymous):

the first part was the product rule.

OpenStudy (anonymous):

2xy+y^2=5 = product rule on (2xy) + differentiate with respect to x (y^2) = 5 except the constant would get removed. so it's actually. 2y + (dy/dx)(2x) + 2y(dy/dx) = 0. sorry.

OpenStudy (anonymous):

ok so would my answer be (-y)/(x+y) when simplified?

OpenStudy (amistre64):

factor out the dy/dx and rearrange for the solve

OpenStudy (amistre64):

-2y/(2x+2y) yeah

OpenStudy (amistre64):

the 2s factor out

OpenStudy (anonymous):

yay! I was hoping I got this one correct

OpenStudy (anonymous):

good job :)